OpenStudy (anonymous):

Complete the square! Ending form 3x^2-6x=2. I added(b/2)^2 to both sides but not really sure where to go from here.

OpenStudy (anonymous):

write 3x^2-6x-2=0

OpenStudy (anonymous):

or write x^2-2x=2/3

OpenStudy (anonymous):

then add 1 on both sides

OpenStudy (anonymous):

u get (x-1)^2=5/3

OpenStudy (anonymous):

and then x-1=square root of (5/3)

OpenStudy (anonymous):

x=1 p;us minus sqroot 5/3

OpenStudy (anonymous):

why do you add 1 to each side?

OpenStudy (anonymous):

in order to complete the formula of completing square on both sides

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

main que is to solve the equation by completing square??

OpenStudy (anonymous):

i had written each thing above

OpenStudy (anonymous):

ok but can't we solve it by quadratic of ehich class it is

OpenStudy (anonymous):

My professor wants it solved by completing the square. It says using any other method will earn no points.

OpenStudy (anonymous):

then do the steps as i have done above but i don't know why i dont get same answers from both methods